KTeV Day Shift, 29-SEP-99

Herman B. White (hwhite@fnal.gov)
Wed, 29 Sep 1999 15:40:07 -0500 (CDT)

Shift Summary for day shift, 28 Sept 99


On Shift: R. Barbosa H. White (also present: Jason H., Jason L.,
Arun T., LEO B.,....

Hours of beam available at all intensities: 7/8.5 hrs
Hours of beam available at nominal intensity: 6/7 hrs
Hours of E799 data (special runs not included): 6 hrs
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beam extraction problems.

========================================================================

Runs ended this shift:

Run 14757 E799 data in progress

=======================================================================

Tapes brought to FCC this shift: OnSplit tapes vaulted:
(Arun T.) UPE002, UPF005-009, UPH002-004, UPK004-005, UPLA01
UPM005-008, UPMU01, UPPE01, UPT002-003, UPX003

==================================================================

Description of shift: We started the shift with RC10/SEM alarms, with
readings about a factor of 10 low, and just beyond alarm limit. We were
informed that the scalar limits were adjusted last night, and some
tuning of the limits is required. The KTeV3a disk were full, and some
files were deleted to address that alarm.
Smooth running most of shift, with available beam.

Occasional beam aborts. Beam off ~1.75hrs QXR extraction problem

===================================================================

run spill time comment
-----------------------------------------------------------------
14757 ~ 09:27 DAQ monitoring display window hangs, unable to
restart.
14757 11:00 minor intensity variation.
11:45 Split tape request KTeV4-tps3d2, UPH005
13:10 QXR problem in the beam spill from MCR.
Request that intensity is lowered to 1 turn
until extraction problem is solved.
14:05 Still Investigating. No estimate for beam
14:52 Qxr problem fixed; beam returns
14:53 KTeVANA errors ; also continue to experience scalar
limit problems, with alarms. Beam position being
ajusted.
15:20 Automatic tuning in the MCR coupled with frequent
higher intensity pulses continue, causing scalar
limit alarms.
15:30 Turn over to next crew for the last 30 mins. of beam


================================================================

Follow up on next shift: Label tapes, current supply ~2 days.